# GITA: Graph to Visual and Textual Integration for Vision-Language Graph Reasoning Project Page: v-graph.github.io

### v4

@@ Abstract
Large Language Models (LLMs) are increasingly used for various tasks with graph structures. Though LLMs can process graph information in a textual format, they overlook the rich vision modality, which is an intuitive way for humans to comprehend structural information and conduct general graph reasoning. The potential benefits and capabilities of representing graph structures as visual images (i.e., visual graph) are still unexplored. To fill the gap, we innovatively propose an end-to-end framework, called Graph to vIsual and Textual IntegrAtion (GITA), which firstly incorporates visual graphs into general graph reasoning. Besides, we establish Graph-based Vision-Language Question Answering (GVLQA) dataset from existing graph data, which is the first vision-language dataset for general graph reasoning purposes. Extensive experiments on the GVLQA dataset and five real-world datasets show that GITA outperforms mainstream LLMs in terms of general graph reasoning capabilities. Moreover, We highlight the effectiveness of the layout augmentation on visual graphs and pretraining on the GVLQA dataset.
@@ 1 Introduction
Graph reasoning tasks are pivotal in domains such as recommendation systems (He et al., 2020; Wang et al., 2021), social network analysis (Cao et al., 2020; Huang et al., 2021), and knowledge graph reasoning (Zhang et al., 2021; Liu et al., 2022). Various architectures have been developed, from shallow embedding methods (Bordes et al., 2013; Socher et al., 2013) to advanced Graph Neural Networks (GNNs) (Kipf & Welling, 2016; Xu et al., 2018) and graph Transformers (Zhang et al., 2020; Kreuzer et al., 2021; Chen et al., 2022). While these models excel in graph reasoning tasks, they often lack generalizability, flexibility, and user-friendliness. Firstly, achieving good performance with these models typically requires domain-specific tuning, which limits their abilities to generalize across different domains. Additionally, these models struggle to handle diverse tasks with the same architecture. Each task often requires a specialized design, including task-specific data processing and decoders, leading to limited flexibility. Lastly, compared to the Large Language Models (LLMs) that can engage in conversations with users, these models are less user-friendly.
@@
In contrast, LLMs have shown exceptional generalization capabilities across a diverse spectrum of reasoning tasks (Wei et al., 2022; Yu et al., 2024; Zhou et al., 2023) by encapsulating task-specific demands within a cohesive and interpretable mechanism - text prompts, under a unified architecture with minimal domain-specific adjustments. Those advantages have sparked investigations into the potential of LLMs for graph reasoning. Recent developments lend credence to the notion that LLMs can indeed interpret and manipulate graph-structured data through textual representations. For example, InstructGLM (Ye et al., 2023), GPT4Graph (Guo et al., 2023), and LLMtoGraph (Liu & Wu, 2023) convert graphs into textual descriptions and then use those descriptions paired with queries to enable LLMs to generate accurate responses for graph reasoning tasks. Furthermore, the introduction of benchmarks such as GraphQA (Fatemi et al., 2024) and NLGraph (Wang et al., 2023) is a testament to the growing interest in evaluating LLMs’ effectiveness on graph reasoning tasks framed in natural languages.
@@
Despite the development of numerous methods and benchmarks for graph reasoning on LLMs, they often overlook the valuable vision modality, which however is a natural means for humans to comprehend structural information and has demonstrated its success in various visual reasoning scenarios Hudson & Manning (2019); Zellers et al. (2019); Zhang et al. (2019); Johnson et al. (2017); Andreas et al. (2016); Suhr et al. (2019). Consequently, the following questions arise: (1) Can incorporating visual information be beneficial for general graph reasoning scenarios? (2) If so, how can we effectively integrate the vision modality into graph reasoning? To the best of our knowledge, these questions remain unexplored.
@@
To answer them, we first propose an end-to-end framework called Graph to vIsual and Textual IntegrAtion (GITA) that systematically integrates visual information into instruction-based graph reasoning, by rendering graph structures to customized visual images which are called visual graph. Specifically, the GITA framework has four components: a graph visualizer for generating visual graphs, a graph describer for producing textual descriptions of the graph structure, a task-based questioner that organizes the description and requirements of the current task into prompt instruction, and a Vision-Language Model (VLM) to perform vision-language graph reasoning. According to the proposed GITA framework, visual information can be incorporated into many tasks with explicit or implicit graph structures, without sacrificing its versatility, flexibility, or user-friendliness. Besides, since there is no dataset for vision-supported general graph reasoning capabilities, we establish the first vision-language dataset for general graph reasoning purposes called Graph-based Vision-Language Question Answering (GVLQA) based on the proposed GITA framework. The GVLQA dataset consists of 526K instances covering seven representative graph reasoning tasks, aiming to thoroughly evaluate the structure-based graph reasoning abilities of VLMs and LLMs. Extensive experiments on the GVLQA dataset and five real-world datasets demonstrate the effectiveness of the proposed GITA model. Furthermore, we delve into the effects of visual graph augmentation strategies and find that layout augmentation can dramatically boost vision-based graph reasoning performance.
@@
  • Extensive experiments on benchmark datasets across various graph reasoning tasks demonstrate the effectiveness of the proposed GITA framework, and highlight the benefits of layout augmentation on visual graphs.
@@ 2 Related Work
Graph Reasoning. Graph reasoning (Battaglia et al., 2018; Wu et al., 2020) aims to answer questions based on graphs, which involves utilizing graph structures to guide the reasoning process to generate answers. Graph reasoning has a wide variety of applications in social network analysis (Newman, 2003; Leskovec et al., 2008), bioinformatics (Jeong et al., 2001; Gavin et al., 2006), chemistry (Gilmer et al., 2017), physics (Battaglia et al., 2016), knowledge graph reasoning (Bordes et al., 2013), and recommendation systems (Koren et al., 2009; He et al., 2017). Many graph reasoning methods have been proposed. Early attempts Bordes et al. (2013); Socher et al. (2013) learn node and edge representations through shallow modules, which may have only limited expressive power. Graph Neural Networks (GNNs) such as GCN Kipf & Welling (2016), GAT Velickovic et al. (2017), GraphSAGE Hamilton et al. (2017), MPNN Gilmer et al. (2017), and GIN Xu et al. (2018) use message-passing paradigm Gilmer et al. (2017) to model graph dependencies and update node features. Transformer-based graph models Zhang et al. (2020); Kreuzer et al. (2021); Chen et al. (2022) further propose to use self-attention to increase the expressiveness and long-range dependency. However, as discussed in Sec 1, these models may exhibit limited generalizability, flexibility, and user-friendliness.
@@
LLMs on Graph Reasoning. There have been many attempts to use LLMs in graph reasoning. Depending on how they align the input spaces of graphs and LLMs, we categorize them into two types: Graph-to-text and Graph-to-token. Graph-to-text methods transform a graph into textual descriptions, which are concatenated with the instructions and fed to the LLM for querying. For example, InstructGLM (Ye et al., 2023) uses natural language to describe the graph and proposes instruction prompts to fine-tune the LLM. He et.al He et al. (2024) applies LLMs to explain graphs for training GNNs, while Chen et.al Chen et al. (2023) treat LLMs as enhancers to exploit text attributes or as predictors for node classification on text-attributed graphs. GPT4Graph (Guo et al., 2023) and LLMtoGraph (Liu & Wu, 2023) convert graphs into specific code or natural language formats by the powerful ChatGPT (OpenAI, 2022, 2023). On the other hand, Graph-to-token methods include GraphGPT (Tang et al., 2023), GraphToken (Perozzi et al., 2024) and LLaGA (Chen et al., 2024). For these methods, the graph is represented as a specially designed token sequence, which is projected or merged into the LLM’s token space for text-based reasoning. However, none of the aforementioned methods represent the graph structure information as images, highlighting the uniqueness of the proposed GITA framework and GVLQA dataset.
@@
Large Vision-Language Models. Large VLMs have significantly expanded the cognitive abilities of LLMs by integrating the vision modality to address vision-language tasks. Many methods have been proposed. Some early explorations like Flamingo Alayrac et al. (2022), CLIP Radford et al. (2021), and BLIP-2 Li et al. (2023) use a visual encoder for processing images and align the visual and textual embeddings. Subsequent models like LLaVA Liu et al. (2023) and MiniGPT-4 Zhu et al. (2023) combine visual and textual inputs in a single LLM for solving multimodal tasks. InstructBlip Dai et al. (2023) proposes an instruction-aware query transformer and trains a vision-language model by instruction tuning. However, despite progress in a wide range of vision-language tasks (Zhang et al., 2024), using visual information in graph reasoning remains overlooked. We take the first step in this field, pushing the boundaries of VLMs in graph reasoning.
@@ 3 GITA: Graph to Visual and Textual Integration
3.1 Preliminary
Graph Reasoning. In traditional graph reasoning settings, models typically rely on two main inputs: i. the graph structure $G=\{C,E\}$, where $C$ and $E$ are the set of vertices and edges, respectively; ii. the task requirement $T$, encompassing specific operations or questions pertaining to the graph. Based on the information provided in $G$ and a specific task requirement $T$, models are expected to output a reasonable answer $A$. On the other hand, in the context of instruction-based graph reasoning methods, it is necessary to convert these inputs into textual form. This transformation facilities graph reasoning within natural language, allowing for improved interpretation and harnessing the formidable reasoning capabilities of large language models.
3.2 Architecture
Overview. Different from the above graph reasoning methods, we propose a Graph to Image-Txt Assistant (GITA), which is the first attempt to perform graph reasoning in a vision-text-based manner. GITA comprises four pivotal components: a task-agnostic graph visualizer $V$, a graph describer $D$, a task-specific questioner $Q$, and a VLM reasoner $R_{\phi}$, as illustrated in Figure 1. Firstly, $V$ and $G$ are designed to produce visual depictions (i.e., visual graphs) and textual descriptions of the graph structure inputs, respectively. Then, given the task requirement $T$ and the textual description produced by $D$, $Q$ is designed to formulate a task-specific query. Finally, $R_{\phi}$ receives the visual graph from $V$ and the task-specific query from $Q$ as the visual input $I_{G}$ and textual input $Q_{G}^{T}$, respectively, consequently generating answers $A$ in natural language.
Graph Visualizer. The role of the graph visualizer is to generate visual graphs from structural graphs. Representing a structural graph with an image is not unique, as there can be variations in many aspects, such as backdrop colors, layouts, and node shapes. These variations may enhance the robustness of models through effective training but simultaneously increase the learning difficulty for models. Therefore, a careful trade-off between consistency and variety is necessary during the graph visualization process. This trade-off is reflected in our design of graph visualizer, by maintaining consistency in basic image styles common to general images (i.e., size, resolution, backdrop, etc.) and only introducing customizable variations in four graph-related image styles unique to visual graphs (i.e., layout, node shapes, node outline styles, and edge thickness). Hence, the graph visualization of $V$ can be concisely encapsulated by the following equation:
@@ 3.2 Architecture
Besides, visualizing the entire graph can be challenging when the number of nodes or edges is very large, affecting the clarity of the images. To address this, our graph visualizer adopts a typical strategy of $k$-hop subgraph sampling. Specifically, $k$-hop subgraph sampling for a node $u$ in the set of vertices $C$ involves selecting a subgraph $G_{u}=\{C_{u}\subseteq N_{k}(u),E_{u}\subseteq E\}$, where $N_{k}(u)$ includes nodes within $k$ steps from $u$ and each edge $(i,j)$ in $E_{u}$ connects nodes within $C_{u}$. To generate the visual graph of the $k$-hop subgraph $G_{u}$ centered on $u$, the nodes within $G_{u}$ will be relabeled from 0 to $|C_{u}|-1$ to facilitate the generalization of visual graphs. Subsequently, this relabeled subgraph $G_{u}$ is input into the graph visualizer to generate its the visual graph $I_{G_{u}}$ as shown in Eq. 1.
In practice, the graph visualizer can be implemented by a variety of graphic visualization tools, such as Graphviz (Gansner & North, 2000), Matplotlib (Tosi, 2009), and NetworkX (Hagberg et al., 2008). Among them, Graphviz can automatically design the layouts of visual graphs, especially suitable for building large-scale datasets. Matplotlib is excellent for customizable plots with fine-grained control, and NetworkX excels in complex network analysis. We have implemented various graph visualizers using modular, plug-in architecture in GITA. Specific examples of the visual graphs generated with these tools can be found in Appendix C.
Graph Describer. The graph describer $D$ is tasked with generating task-agnostic textual descriptions of a given graph $G$. To ensure the clarity and fidelity of these descriptions, we meticulously craft a curated set of graph-describing templates. The graph description templates outlined in Appendix D are designed to cover a broad spectrum of scenarios, accommodating various graph configurations including directed or undirected graphs and those with or without node or edge weights. To generate a specific description for a given graph, the graph describer initially selects the appropriate template based on the graph’s characteristics, such as its directionality and whether it includes node attributes or edge weights. Subsequently, this template is used by replacing placeholders with actual data, such as the number of nodes, the number of edges, and the endpoints of each edge, to craft detailed descriptions tailored to the specific graph in question. The process for $D$ to generate textual descriptions can be formulated as follows:
@@
Questioner. The questioner $Q$, is tailored to capture the intricate requirements of specific tasks and reflect them in its output task-specific query. In detail, $Q$ receives the task-agnostic textual descriptions from the graph describer and refines them to align with the task context by elucidating the concrete meanings of nodes and edges. These refined descriptions are then enriched with task responsibilities and input/output specifications to form task-specific queries. The formulation of the questioner to generate the task-specific queries can be represented as follows:
@@
where $Q_{G}^{T}$ represents the task-specific query generated by the questioner with given the task requirement $T$ and the textual description $D_{G}$. The construction of task-specific queries can be approached in two main ways: manual-template-based constructing and bootstrapping LLM agents. manual-template-based constructing enriches $D_{G}$ with task-specific manual templates, which is preferred for tasks with precise requirements, such as the Traveling Salesman Problem (TSP) (Dantzig et al., 1954), where accuracy is critical and the task definitions are well-understood. This is because it can ensure clarity and reduces the risk of errors due to its meticulous attention to detail. On the other hand, bootstrapping LLM agents for automated synthesis is more economic and suitable for dynamic or bespoke tasks, such as robotic planning or complex gaming scenarios, as it can take advantage of the speed and adaptability of LLM agents to interpret context and generate appropriate queries, minimizing manual effort and enhancing responsiveness to changing conditions. Both methods are illustrated with examples in Appendix E, showcasing their applications and benefits in different scenarios.
VLM Reasoner. The VLM reasoner $R_{\phi}$ is to perform the final graph reasoning with visual inputs $I_{G}$ from $V$ and textual inputs $Q_{G}^{T}$ from $Q$, consequently outputting responses in natural languages. This reasoning process can be presented as the following formula:
@@
where $A$ is the answer generated by a vision-language model $R$. In this work, we specifically adopt GPT-4V and LLaVA-7B/13B as VLM reasoners. Those models are regarded as representatives in the realm of closed-source and open-source VLMs, respectively.
In summary, GITA systematically incorporates the vision modality into instruction-based graph reasoning. In Appendix A, we discuss the characteristics of GITA, in aspects of generalizability, flexibility, user-friendliness, and large-graph scalability.

### v5

@@ Abstract
Large Language Models (LLMs) are increasingly used for various tasks with graph structures. Though LLMs can process graph information in the textual format, they overlook the rich vision modality, which is an intuitive way for humans to comprehend structural information and conduct general graph reasoning. The potential benefits and capabilities of representing graph structures as visual images (i.e., visual graph) are still unexplored. To fill the gap, we innovatively propose an end-to-end framework, called Graph to vIsual and Textual IntegrAtion (GITA), which incorporates visual graphs into general graph reasoning. Besides, we construct the Graph-based Vision-Language Question Answering (GVLQA) dataset from existing graph data, which is the first vision-language dataset for general graph reasoning. Extensive experiments on the GVLQA dataset and five real-world datasets show that GITA outperforms mainstream LLMs on general graph reasoning. Moreover, experimental results demonstrate the effectiveness of the layout augmentation on visual graphs and pretraining on the GVLQA dataset.
@@ 1 Introduction
Graph reasoning tasks are pivotal in domains such as recommendation systems [25; 60], social network analysis [7; 29], and knowledge graph reasoning [72; 46; 62]. Various architectures have been developed, from shallow embedding methods [6; 53] to advanced Graph Neural Networks (GNNs) [37; 64] and graph Transformers [71; 40; 8]. While these models excel in graph reasoning tasks, they often lack generalizability, flexibility, and user-friendliness. Achieving good performance with these models typically requires domain-specific tuning, which limits their abilities to generalize across different domains. Additionally, these models struggle to handle diverse tasks with the same architecture. Each task often requires a specialized design, including task-specific data processing and decoder, leading to limited flexibility. Lastly, unlike the Large Language Models (LLMs) that can engage in conversations with users, these models are less explainable and user-friendly.
@@
In contrast, LLMs have shown great generalization capabilities across a wide variety of reasoning tasks [61; 67; 74; 32; 33] by encapsulating task-specific demands within a cohesive and interpretable mechanism - text prompts, under a unified architecture with minimal domain-specific adjustments. These advantages have sparked investigations into the potential of LLMs for graph reasoning. Recent developments lend credence to the notion that LLMs can indeed interpret and manipulate graph-structured data through textual representations. For example, InstructGLM [66], GPT4Graph [21], and LLMtoGraph [44] convert graphs into textual descriptions and then use these descriptions paired with queries to enable LLMs to generate accurate responses for graph reasoning tasks. Furthermore, the introduction of benchmarks such as GraphQA [14] and NLGraph [59] is a testament to the growing interest in evaluating LLMs’ effectiveness on graph reasoning tasks framed in natural languages.
@@
Despite the development of numerous methods and benchmarks for graph reasoning on LLMs, they often overlook the valuable vision modality, which is a natural means for humans to comprehend structural information and has demonstrated its success in various visual reasoning scenarios [30; 68; 69; 34; 3; 54]. Consequently, the following questions arise: (1) Can incorporating visual information be beneficial for general graph reasoning scenarios? (2) If so, how can we effectively integrate the vision modality into graph reasoning? To the best of our knowledge, these questions remain unexplored.
@@
To answer them, we first propose an end-to-end framework called Graph to vIsual and Textual IntegrAtion (GITA) [^1] Project Homepage: v-graph.github.io. [^2] Code Repository: https://github.com/WEIYanbin1999/GITA/. that systematically integrates visual information into instruction-based graph reasoning, by rendering graph structures to customized visual images which are called visual graph. Specifically, the GITA framework has four components: a graph visualizer for generating visual graphs, a graph describer for producing textual descriptions of the graph structure, a task-based questioner that organizes the description and requirements of the current task into prompt instruction, and a Vision-Language Model (VLM) to perform vision-language graph reasoning. In the proposed GITA framework, visual information can be incorporated into many tasks with explicit or implicit graph structures, without sacrificing its versatility, flexibility, or user-friendliness. Besides, since there is no dataset for vision-supported general graph reasoning capabilities, we construct the first vision-language dataset for general graph reasoning purposes called Graph-based Vision-Language Question Answering (GVLQA) [^3] Dataset: https://huggingface.co/collections/Yanbin99/. based on the proposed GITA framework. The GVLQA dataset consists of 526K instances covering seven representative graph reasoning tasks, aiming to thoroughly evaluate the structure-based graph reasoning abilities of VLMs and LLMs. Extensive experiments on the GVLQA dataset and five real-world datasets demonstrate the effectiveness of the proposed GITA model. Furthermore, we delve into the effects of visual graph augmentation strategies and find that layout augmentation can dramatically boost vision-based graph reasoning performance.
@@
  • Extensive experiments on benchmark datasets across various graph reasoning tasks demonstrate the effectiveness of the proposed GITA framework and the benefits of layout augmentation on visual graphs.
@@ 2 Related Work
Graph Reasoning. Graph reasoning [5; 63] aims to answer questions based on graphs, which involves utilizing graph structures to guide the reasoning process to generate answers. Graph reasoning has a wide variety of applications in social network analysis [47; 41], bioinformatics [31; 18], chemistry [19], physics [4], knowledge graph reasoning [6], and recommendation systems [39; 26]. Many graph reasoning methods have been proposed. Early attempts [6; 53] learn node and edge representations through shallow modules, which may have only limited expressive power. Graph Neural Networks (GNNs) such as GCN [37], GAT [58], GraphSAGE [23], MPNN [19], and GIN [64] use message-passing paradigm [19] to model graph dependencies and update node features. Transformer-based graph models [71; 40; 8] further propose to use self-attention to increase the expressiveness and long-range dependency. However, as discussed in Sec 1, these models may exhibit limited generalizability, flexibility, and user-friendliness.
@@
LLMs on Graph Reasoning. There have been many attempts to use LLMs in graph reasoning. Depending on how they align the input spaces of graphs and LLMs, we categorize them into two types: Graph-to-text and Graph-to-token. Graph-to-text methods transform a graph into textual descriptions, which are concatenated with the instructions and fed to the LLM for querying. For example, InstructGLM [66] uses natural language to describe the graph and proposes instruction prompts to fine-tune the LLM. He et.al [27] applies LLMs to explain graphs for training GNNs, while Chen et.al [10] treat LLMs as enhancers to exploit text attributes or as predictors for node classification on text-attributed graphs. GPT4Graph [21] and LLMtoGraph [44] convert graphs into specific code or natural language formats by the powerful ChatGPT [48; 49]. On the other hand, Graph-to-token methods include GraphGPT [55], GraphToken [50] and LLaGA [9]. For these methods, the graph is represented as a specially designed token sequence, which is projected or merged into the LLM’s token space for text-based reasoning. However, none of the aforementioned methods represent the graph structure information as images, highlighting the uniqueness of the proposed GITA framework and GVLQA dataset.
@@
Large Vision-Language Models. Large VLMs have significantly expanded the cognitive abilities of LLMs by integrating the vision modality to address vision-language tasks. Many methods have been proposed. Some early explorations like Flamingo [2], CLIP [51], and BLIP-2 [43] use a visual encoder for processing images and align the visual and textual embeddings. Subsequent models like LLaVA [45] and MiniGPT-4 [75] combine visual and textual inputs in a single LLM for solving multimodal tasks. InstructBlip [11] proposes an instruction-aware query transformer and trains a vision-language model by instruction tuning. However, despite progress in a wide range of vision-language tasks [70; 16], using visual information in graph reasoning remains overlooked. We take the first step in this field, pushing the boundaries of VLMs in graph reasoning.
@@ 3 GITA: Graph to Visual and Textual Integration
3.1 Preliminary
Graph Reasoning. In traditional graph reasoning settings, models typically rely on two main inputs: (i) the graph structure $G=\{C,E\}$, where $C$ and $E$ are the set of vertices and edges, respectively; (ii) the task requirement $T$, encompassing specific operations or questions pertaining to the graph. Based on the information provided in $G$ and a specific task requirement $T$, models are expected to output a reasonable answer $A$. On the other hand, in the context of instruction-based graph reasoning methods, it is necessary to convert these inputs into textual form. This transformation facilities graph reasoning within natural language, allowing for improved interpretation and harnessing the formidable reasoning capabilities of large language models.
3.2 Architecture
Overview. Different from the above graph reasoning methods, we propose a Graph to Image-Txt Assistant (GITA), which is the first attempt to perform graph reasoning in a vision-text-based manner. GITA comprises four pivotal components: a task-agnostic graph visualizer $V$, a graph describer $D$, a task-specific questioner $Q$, and a VLM reasoner $R_{\phi}$, as illustrated in Figure 1. Firstly, $V$ and $G$ are designed to produce visual depictions (i.e., visual graphs) and textual descriptions of the graph structure inputs, respectively. Then, given the task requirement $T$ and the textual description produced by $D$, $Q$ is designed to form a task-specific query. Finally, $R_{\phi}$ receives the visual input $I_{G}$ from $V$ based on the visual graph and the textual input $Q_{G}^{T}$ from $Q$, then generates answers $A$ in natural language. In the following, we introduce the four components in detail.
@@ 3.2 Architecture
Graph Visualizer. The role of the graph visualizer is to generate visual graphs from structural graphs. The image representation of a structural graph is not unique, as there can be variations in many aspects, such as backdrop colors, layouts, and node shapes. These variations may enhance the robustness of models through effective training but simultaneously increase the learning difficulty for models. Therefore, balancing consistency and variety is necessary during the graph visualization process. This trade-off is reflected in our design of graph visualizer, by maintaining consistency in basic image styles common to general images (i.e., size, resolution, backdrop) and only introducing customizable variations in four graph-related image styles unique to visual graphs (i.e., layout, node shapes, node outline styles, and edge thickness). Graph visualization in $V$ can be formulated by the following equation:
@@
Visualizing the entire graph can be challenging when the number of nodes or edges is very large, affecting the clarity of the images. To address this, our graph visualizer adopts the standard strategy of $k$-hop subgraph sampling. Specifically, $k$-hop subgraph sampling for a node $u$ in the set of vertices $C$ involves selecting a subgraph $G_{u}=\{C_{u}\subseteq N_{k}(u),E_{u}\subseteq E\}$, where $N_{k}(u)$ includes nodes within $k$ steps from $u$ and each edge $(i,j)$ in $E_{u}$ connects nodes within $C_{u}$. To generate the visual graph of the $k$-hop subgraph $G_{u}$ centered on $u$, the nodes within $G_{u}$ are relabeled from 0 to $|C_{u}|-1$ to facilitate the generalization of visual graphs. Subsequently, this relabeled subgraph $G_{u}$ is fed to the graph visualizer to generate its visual graph $I_{G_{u}}$ by Eq. (1).
In practice, the graph visualizer can be implemented by a variety of graphic visualization tools, such as Graphviz [17], Matplotlib [56], and NetworkX [22]. Among them, Graphviz can automatically design the layouts of visual graphs, and is especially suitable for building large-scale datasets. Matplotlib is excellent for customizable plots with fine-grained control, and NetworkX excels in complex network analysis. We have implemented various graph visualizers using modular, plug-in architecture in GITA. Specific examples of the visual graphs generated with these tools can be found in Appendix D.
Graph Describer. The graph describer $D$ is tasked with generating task-agnostic textual descriptions of a given graph $G$. To ensure clarity and fidelity of these descriptions, we meticulously craft a curated set of graph-describing templates. The graph description templates outlined in Appendix E are designed to cover a broad spectrum of scenarios, accommodating various graph configurations including directed or undirected graphs and those with or without node or edge weights. To generate the description for a given graph, the graph describer initially selects an appropriate template based on the graph’s characteristics, such as its directionality and whether it includes node attributes or edge weights. Subsequently, this template is used by replacing placeholders with actual data, such as the number of nodes, the number of edges, and the endpoints of each edge, to craft detailed descriptions tailored to the specific graph in question. The process for $D$ to generate textual descriptions can be formulated as follows:
@@
Questioner. The questioner $Q$ is tailored to capture the intricate requirements of specific tasks and reflect them in its output task-specific query. In detail, $Q$ receives the task-agnostic textual descriptions from the graph describer and refines them to align with the task context by elucidating the concrete meanings of nodes and edges. These refined descriptions are then enriched with task responsibilities and input/output specifications to form task-specific queries. The formulation of the questioner to generate the task-specific queries can be represented as follows:
@@
where $Q_{G}^{T}$ represents the task-specific query generated by the questioner with given the task requirement $T$ and the textual description $D_{G}$. The construction of task-specific queries can be approached in two main ways: manual template-based construction and bootstrapping LLM agents. Manual template-based construction enriches $D_{G}$ with task-specific manual templates, which is preferred for tasks with precise requirements, such as the Traveling Salesman Problem (TSP) [12], where accuracy is critical and the task definitions are well-understood. This is because it can ensure clarity and reduce the risk of errors due to its meticulous attention to details. On the other hand, bootstrapping LLM agents for automated synthesis is more economical and suitable for dynamic or bespoke tasks, such as robotic planning or complex gaming scenarios, as it can take advantage of the speed and adaptability of LLM agents to interpret context and generate appropriate queries, minimizing manual effort and enhancing responsiveness to changing conditions. Both methods are illustrated with examples in Appendix F, showcasing their applications and benefits in different scenarios.
VLM Reasoner. The VLM reasoner $R_{\phi}$ performs final graph reasoning with visual inputs $I_{G}$ from $V$ and textual inputs $Q_{G}^{T}$ from $Q$, and outputs responses in natural language. This reasoning process can be represented as the following:
@@
where $A$ is the answer generated by the vision-language model $R$. In this work, we adopt GPT-4V and LLaVA-7B/13B as VLM reasoners. These models are regarded as representatives in the realm of closed-source and open-source VLMs, respectively.
In summary, GITA systematically incorporates the vision modality into instruction-based graph reasoning. In Appendix B, we discuss the characteristics of GITA, in aspects of generalizability, flexibility and user-friendliness.